OpenStudy (carraway28):

simplify 1times 0 - 0 over y

OpenStudy (phi):

do you mean \[ 1 \times 0 - \frac{0}{y} \]

OpenStudy (carraway28):

yea

OpenStudy (carraway28):

like how would i work that out?

OpenStudy (phi):

do you know order of operations ?

OpenStudy (carraway28):

im havin a hard time rememberin bc this was from 9th grade an its been awhile im have to make it up bc i didnt do good enough on it

OpenStudy (phi):

I've seen people use PEDMAS Parentheses (we don't have any in your problem) Exponents (like \( 3^2 \) for example. we don't have those either) Divide (we have that) Multiply (we have that) Add (we have this) Subtract (we don't have this)

OpenStudy (phi):

so you should do the divide first. that is 0/y any idea what that is ?

OpenStudy (carraway28):

how can you divide 0/y??

OpenStudy (phi):

it helps to know it is the same as \[ 0 \times \frac{1}{y} \] and there is a "rule" zero times anything is what ?

OpenStudy (carraway28):

0

OpenStudy (phi):

yes, so 0 * 1/y or 0/y is 0

OpenStudy (carraway28):

ok then it would b 1x0 -0

OpenStudy (phi):

yes. Next we do the Multiply 1*0 is ?

OpenStudy (carraway28):

0

OpenStudy (phi):

yes now you have 0 - 0 that is a Subtract, and the last operation we do.

OpenStudy (carraway28):

so the answer would b 0.....thank you for helping me

OpenStudy (phi):

yes. the only thing you should be careful of, is if y was 0 then we would have 0/0 and that is not defined so the "complete answer" would be \[ 0 \ \text{ if } y \ne 0\\ \text{undefined if } \ y =0 \]

OpenStudy (carraway28):

o ok thank you